OpenStudy (vishweshshrimali5):

find the locus of the point P suchh that |PA - PB| = 6 where A = (0,4) and B = (0,-4).

OpenStudy (vishweshshrimali5):

@hartnn

OpenStudy (vishweshshrimali5):

I tried in this way \[|\sqrt{h^2 + (k-4)^2} - \sqrt{h^2 + (k+4)^2}| = 6\]

OpenStudy (vishweshshrimali5):

Then I squared both sides but couldn't solve further

OpenStudy (vishweshshrimali5):

HERE (h,k) = P

hartnn (hartnn):

thats what came to my mind first , and i requires lot of algebra... when u square it, u get one more sqrt(...) term, which u need t0 isolate, keep it on one side of = sign and the square again

hartnn (hartnn):

very tedious algebra,

OpenStudy (vishweshshrimali5):

@estudier @calculusfunctions

OpenStudy (anonymous):

take x=cosP y-4=sinP Hope it will make your equation a lot easier

OpenStudy (vishweshshrimali5):

What is P here

OpenStudy (vishweshshrimali5):

@akkib89

OpenStudy (anonymous):

P is tan inverse (y-4)/x

OpenStudy (vishweshshrimali5):

Can u help by starting with 2 or 3 steps

OpenStudy (anonymous):